OpenStudy (anonymous):

How do you solve: A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 12 + 42 + 72 + . . . + (3n - 2)2 = (n(6n^2-3n-1))/2

OpenStudy (anonymous):

replace \(n\) by 1 to get \[S_1\] which will be \[12=\frac{1(6\times 1^2-3\times 1-1)}{2}\]which does not seem to be true does it?

OpenStudy (anonymous):

no it does not it equals 0

OpenStudy (anonymous):

and sorry it was 1^2+4^2+7^2 then (3n-2)^2